OpenStudy (anonymous):

If the domain of f(x) = (1 / 2^x) - 7 is all integers, what is a two digit positive integer that is in the range of f(x)?

OpenStudy (anonymous):

is it \(\frac{1}{2^x}-7\)?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

try negative values for \(x\) not that hard since \(\frac{1}{2^x}=x^{-1}\)

OpenStudy (anonymous):

you could take say \(x=-4\) and get \(2^4-7=16-7=9\) but that is a one digit number, so not an answer try \(x=-5\)

OpenStudy (anonymous):

ohh i see thanks!

OpenStudy (anonymous):

yw